# Discovering Artificial Viscosity Models for Discontinuous Galerkin approximation of Conservation Laws using Physics-Informed Machine Learning

### v1

@@ 4.3 Training in practice
In this section, we provide more details on how the training is carried out in the one-dimensional ($d=1$) case. This procedure can be easily generalized to different dimensions. The first step is to choose which problems to train the neural network on, namely, we need to choose different combinations $\{\mathscr{P}_{j}\}_{j=1}^{N_{p}},\{\mathscr{P}_{j}^{\text{test}}\}_{j=1}^{N_{pt}}$ of problems on which to train and test. The two sets are an 80-20 random split of a union of cartesian products of parameter spaces. For the 1D case we consider uniform discretizations of $(0,1)$ with $h=\frac{2^{-i}}{10},i=1,...,4$, degrees $k=1,2,4$, the flux $\boldsymbol{f}$ of linear advection, Burgers’ equation and Euler system, a constant timestep $\Delta t$ that is the largest such that the considered problem is stable (it varies from problem to problem), number of timesteps $N=100$ for linear advection and appositely chosen as the smallest $N\in\mathbb{N}$ such that shocks and waves appear in Burgers’ and Euler cases. Concerning initial conditions, we use the functions reported on the left of Table 1 for advection and Burgers’ problems, and a Sod shock tube problem [57] for Euler’s. A representation of the initial conditions is shown in Figure 9. For the 2D case, the problems are chosen in an analogous way apart from the following parameters. Since unstructured grids offer flexibility in discretizing computational domains, allowing for efficient representation of intricate shapes and unbounded regions, we use both structured and unstructured triangulations $\mathcal{K}_{h}$ to train the NN. Namely, we discretize $(0,1)^{2}$ with $h=\frac{2^{-i}}{10},i=1,...,3$. The initial conditions are shown on the right of Table 1 for advection and Burgers’ equations. We also add the Euler’s system in the Riemann configuration 4 of [36] to the training environment. Moreover, in all the considered problems the exact solution $\boldsymbol{u}_{ex}$ is either the analytical solution (when available) or an overkill solution obtained using the DG solver on a triangulation that has a mesh size $h$ that is at least eight times smaller.
@@
  • the jump across the element, the gradient of the velocity, and the velocity at the previous timestep are all features that significantly help the training;
  • the Gaussian error linear unit (GeLU) [29] is the activation function that performed better when compared to the rectified linear unit (ReLU), exponential linear unit (ELU), sigmoid linear unit (SiLU) and hyperbolic tangent (Tanh);
  • a large model with a width greater than 100 seems to perform better, in particular, we choose to use 120 neurons per layer;
@@ 5.1.2 Test case 2: two-dimensional smooth problem
The problem is endowed with periodic boundary conditions. Concerning the discretization, we employ a structured triangular grid of the domain $\Omega=(0,1)^{2}$ of granularity $h=\frac{\sqrt{2}}{2^{i}10},i=1,...,5$. In Figure 7 we report the computed errors in the $L^{2}(\Omega)$-norm and the expected rates of convergence. The NN technique achieves the optimal convergence rate. Numerical experiments show also that we obtain the same roc independently of the mesh type (structured or unstructured).
@@ 5.2 One-dimensional problems
Figure 8: Examples of initial conditions $u_{0}$ used to train the NN.
@@ 5.2.1 Test case 3: linear advection
The first test case we consider is a linear advection problem with a highly discontinuous initial condition, namely (22) in the $\Omega=(0,1)$ domain. The simulation is performed until $T=0.4$ with polynomial degrees $k=1,3,5$ and endowed with periodic boundary conditions. To compare solutions with a similar number of degrees of freedom we select $h=\frac{1}{60},\frac{1}{30},\frac{1}{15}$ and $\text{CFL}=0.2,0.5,0.75$, respectively. The solution obtained with the neural viscosity model is compared with a tuned EV model ($c_{K}=0.6$, $c_{\max{}}=0.3$, see Eq. (11), (12)) and two instances of EV model that are not optimally tuned, namely one case where the injected dissipation is too small and too large. In Figure 10 we compare the solutions at the final time: we can see that the NN based solution is symmetric, essentially non-oscillatory, and well captures the front of the wave. By analyzing the space-time plot of the viscosity introduced via the NN, as shown in Figure 11, we notice that the NN tends to add a larger amount of more localized viscosity when compared with EV. More quantitative results about the error are shown in Table 2, we notice that the NN always outperforms the EV model in terms of $L^{1}$ error and overshoots/undershoots. It also obtains comparable results when considering mass variation (mv), an important metric to asses that the physics of the problem is respected.
@@ 5.2.2 Test case 4: Burgers’ equation
In this test case, we keep the same parameters of the previous test case but we change the physical flux, namely, we employ a quadratic one, and the $\text{CFL}=0.15,0.4,0.4$. These settings are more challenging due to the fact that as the solution evolves rarefaction fans and shock waves develop and collide. We compare the NN solution with a tuned EV model ($c_{K}=3.0,c_{\max{}}=1.0$) and two other not optimally tuned EV models. Figure 12 shows that one of the EV models ($c_{K}=1.0,c_{\max{}}=0.5$) is not able to add enough viscosity thus leading to large oscillation. The parameters used in this case are not far from the manually tuned: this showcases how time-consuming the parameter selection can be. In this case, the NN model is by far the best in avoiding overshoots/undershoots while correctly predicting the profile of the wave. The history of the viscosity in Figure 13 shows that as in the previous test cases, the NN tends to inject more viscosity than the EV model, but in a more localized and less continuous way. For $k=5$ in particular, it is possible to see that the NN also tends to add dissipation in $x=0.2,t=0.35$ where the EV model did not. It remains to be investigated if this is a generalization error (since $k=5$ was not present in the training problems) or a feature of the discovered neural viscosity model. As shown in Table 3, the NN model outperforms all the EV models under all the considered metrics. This is an especially interesting result when considering that the reduction of the $L^{1}$ error for $k=3$ is of 25%, while also reducing overshoots/undershoots of almost 50%.
@@ 5.2.3 Test case 5: Sod shock tube
NN EV NN EV NN EV
$\epsilon$ 5.7046e+01 5.8845e+01 5.8787e+01 7.2200e+01 4.3138e+01 5.7406e+01
$\nabla\epsilon$ 2.7103e+01 2.7001e+01 1.3087e+02 1.3356e+02 2.0463e+02 2.1445e+02
$[\![\epsilon]\!]$ 1.5025e+00 1.8547e+00 1.0929e+00 1.4114e+00 8.7019e-01 1.0149e+00
o/u 3.0200e-01 7.5084e-01 2.0631e-01 7.5516e-01 8.7122e-01 1.1302e+00
@@
It is endowed with constant Dirichlet boundary conditions (cfr. [57]). For the discretization we select three cases with $k=1,3,5$, $h=\frac{1}{60},\frac{1}{30},\frac{1}{15}$, $\text{CFL}=0.27,0.61,0.88$, respectively. This is done to compare solutions with a similar number of degrees of freedom. The discontinuities in the initial conditions give rise to three characteristic waves, demanding robust numerical schemes to accurately compute the solution in the presence of shock waves and rarefactions. Namely, these are a right-moving contact wave and shock wave, and a left-moving rarefaction wave. Classical models aim to add a small amount of viscosity near the contact while constantly introducing dissipation in the region close to the shock. In Figure 14 we compare the NN solution at final time with the EV tune model with $c_{K}=1.0,c_{\max{}}=0.5$. The neural viscosity model provides a solution that has much smaller overshoots (see $x=0.2,k=1$ for instance) and that captures with more precision the wavefront (see $x=0.7,k=1$ for instance). In this case, it is particularly interesting to analyze the history of the viscosity shown in Figure 15. Namely, here we can notice a large discrepancy between what the NN and the EV model aim to do: the EV model stabilizes only the shock, meanwhile the NN wants to stabilize all three present waves. This is counterbalanced by injecting a smaller dissipation near the shock. Despite being more intrusive, the NN outperforms the EV model under all the considered metrics, as shown in Table 4, where we report the error metrics in time and their cumulative value.
@@ 5.2.4 Test case 6: Shu-Osher problem
NN EV NN EV NN EV
$\epsilon$ 2.6293e+03 3.0574e+03 2.2953e+03 4.0510e+03 3.7385e+03 5.5862e+03
$\nabla\epsilon$ 1.7160e+03 1.8731e+03 5.6320e+03 7.8568e+03 1.3576e+04 1.6539e+04
$[\![\epsilon]\!]$ 1.2497e+02 1.0530e+02 9.8602e+01 7.1604e+01 5.2003e+01 5.4259e+01
o/u 3.1512e+00 3.9106e+00 2.8027e+00 5.7163e+00 3.8023e+00 4.6424e+00
@@
The problem is completed with Dirichlet boundary conditions on the left and Neumann on the right part of the boundary (cfr. [56]). Since the solution exhibits a high frequency wave we employ a finer discretization, namely $h=\frac{1}{150},\frac{1}{75},\frac{1}{50}$ with $k=1,3,5$, and $\text{CFL}=0.12,0.3,0.4$, respectively. From Figure 16, we can see that for the case $k=1$ the NN and the EV model produce a rather similar solution. However, the quantitative analysis reported in Table 5 shows that the neural viscosity model is slightly more accurate in the $L^{1}$ norm, and when considering overshoots and undershoots. Instead, the cases $k=3,5$ show that the EV model smooths the solution too much, and thus it is not able to capture the high frequencies. On the other hand, the NN shows much better results, namely the solution correctly represents the high-frequency wave pack present between $x=0.5$ and $x=2.5$. This also proves that the neural viscosity model is able to exploit the high-order nature of the method, indeed the three considered cases have the same number of degrees of freedom. In particular, for $k=3$ we have a reduction of the $L^{1}$ error of 43% and of 51% of overshoot/undershoot. The dissipation injected by the NN, as shown in Figure 17, is concentrated on the shock and, differently from the EV model, adds also other small patches of viscosity at the boundary of the sinusoidal wave pack.
@@ 5.3.1 Test case 7: linear advection
Moreover, it gives us the possibility to check the mass conservation properties of the neural viscosity model. We solve the problem in $\Omega=(0,2)^{2}$ with a final time $T=0.25$ on a structured grid with $7200,1800,450$ triangular elements and $k=1,3,5$, $\text{CFL}=0.075,0.3,0.32$, respectively. We focus on $k=1$ since the other two cases are analogous. As done in the previous test cases, we compare the neural viscosity model with a tuned EV model ($c_{K}=1.0,c_{\max{}}=0.5$) and two other EV models that inject a sub-optimal amount of dissipation (too large $c_{K}=2.0,c_{\max{}}=1.0$ and too small $c_{K}=0.5,c_{\max{}}=0.25$). From Figure 18 we clearly see that both the NN and the EV model smooth the discontinuous profile of the initial condition in order to avoid oscillations. The NN solution is symmetric and visually exhibits a sharper wavefront. This is confirmed by the quantitative analysis of the error reported in Table 6. In particular, it is interesting to observe that the NN model injects more viscosity than a classical EV model but has a smaller $L^{1}$ error. This is not an intuitive behavior when comes to the EV model since from both Figure 19 and Table 6 it is clearly observable that EV models that inject larger dissipation have smaller overshoots/undershoots but larger $L^{1}$ error. Moreover, the solution obtained with the NN conserves the mass better than the one obtained with the EV model, showing that also physics is respected. Therefore, in this case, the neural viscosity model is effectively learning a new model that has better properties than the EV one.
@@ 5.3.2 Test case 8: KPP rotating wave
NN EV NN EV NN EV
$\epsilon$ 1.6196e+05 3.9369e+05 1.6956e+05 3.8346e+05 2.5583e+05 2.8887e+05
$\nabla\epsilon$ 2.4551e+05 4.3755e+05 1.2625e+06 1.7127e+06 2.6152e+06 2.8368e+06
$[\![\epsilon]\!]$ 1.0973e+05 1.4006e+05 8.2736e+04 9.7502e+04 4.4259e+04 6.2624e+04
o/u 8.0091e+03 1.9289e+04 6.4779e+03 6.8667e+03 4.6155e+03 5.0963e+03
@@
The problem is endowed with periodic boundary conditions. For discretization, we employ a structured triangular mesh with $7200,1800,450$ elements and $k=1,3,5$, $\text{CFL}=0.11,0.4,0.48$, respectively. We compare the neural viscosity model with EV models tuned like in the previous test case. The results reported in Figure 2021 and show that the neural viscosity model produces more accurate solutions with respect to the EV one. In particular, it presents less wriggles and a sharper wave front. To understand the behavior of the model we report also in Figure 22 the maximum, mean, and median of the injected viscosity. As was the case in 1D, the NN model injects more viscosity than the tuned EV model, but retains a lower $L^{1}$ error, as shown in Table 7.

### v2

@@ 4.3 Training in practice
In this section, we provide more details on how the training is carried out in the one-dimensional ($d=1$) case. This procedure can be easily generalized to different dimensions. The first step is to choose which problems to train the neural network on, namely, we need to choose different combinations $\{\mathscr{P}_{j}\}_{j=1}^{N_{p}},\{\mathscr{P}_{j}^{\text{test}}\}_{j=1}^{N_{pt}}$ of problems on which to train and test. The two sets are an 80-20 random split of a union of cartesian products of parameter spaces. For the 1D case we consider uniform discretizations of $(0,1)$ with $h=\frac{2^{-i}}{10},i=1,...,4$, degrees $k=1,2,4$, the flux $\boldsymbol{f}$ of linear advection, Burgers’ equation and Euler system, a constant timestep $\Delta t$ that is the largest such that the considered problem is stable (it varies from problem to problem), number of timesteps $N=100$ for linear advection and appositely chosen as the smallest $N\in\mathbb{N}$ such that shocks and waves appear in Burgers’ and Euler cases. Concerning initial conditions, we use the functions reported on the left of Table 1 for advection and Burgers’ problems, and a Sod shock tube problem [57] for Euler’s. A representation of the initial conditions is shown in Figure 9. For the 2D case, the problems are chosen in an analogous way apart from the following parameters. Since unstructured grids offer flexibility in discretizing computational domains, allowing for efficient representation of intricate shapes and unbounded regions, we use both structured and unstructured triangulations $\mathcal{K}_{h}$ to train the NN. Namely, we discretize $(0,1)^{2}$ with $h=\frac{2^{-i}}{10},i=1,...,3$. The initial conditions are shown on the right of Table 1 for advection and Burgers’ equations. We also add the Euler’s system in the Riemann configuration 12 of [36] to the training environment.
@@
  • the gradient of the state variable, the jump across the element, and the state variable at the previous timestep are all features that significantly help the training;
  • the Gaussian error linear unit (GeLU) [29] is the activation function that we choose since it reached the smallest loss when compared to the rectified linear unit (ReLU), exponential linear unit (ELU), sigmoid linear unit (SiLU) and hyperbolic tangent (Tanh);
  • a large model, with a width greater than 100, has a smaller loss on average. In particular, we choose to use 120 neurons per layer;
@@
Therefore, the input vector of the neural network has size 19 (4 for the reference variable, its gradient, its value at the previous time step, its flux; 2 for the jump; 1 for the polynomial degree) in 1D and 29 in 2D (4 for the reference variable, its gradient, its value at the previous time step, its flux and its jump; 1 for the polynomial degree)
@@ 5.1.2 Test case 2: two-dimensional smooth problem
The problem is endowed with periodic boundary conditions. Concerning the discretization, we employ a structured triangular grid of the domain $\Omega=(0,1)^{2}$ of granularity $h=\frac{\sqrt{2}}{2^{i}10},i=1,...,5$. In Figure 7 we report the computed errors in the $L^{2}(\Omega)$-norm and the expected rates of convergence. The NN technique achieves the optimal convergence rate. Numerical experiments show also that we obtain the same rate of convergence independently of the mesh type (structured or unstructured).
@@ 5.2.1 Test case 3: linear advection
The first test case we consider is a linear advection problem with a highly discontinuous initial condition, namely (23) in the $\Omega=(0,1)$ domain. The simulation is performed until $T=0.4$ with polynomial degrees $k=1,3,5$ and endowed with periodic boundary conditions. To compare solutions with a similar number of degrees of freedom we select $h=\frac{1}{60},\frac{1}{30},\frac{1}{15}$, $\text{CFL}=0.2,0.5,0.75$ and $C=0.2,0.0556,0.03$, respectively. The solution obtained with the neural viscosity model is compared with a tuned EV model ($c_{K}=0.6$, $c_{\max{}}=0.3$, see Eq. (12), (13)) and two instances of EV model that are not optimally tuned, namely one case where the injected dissipation is too small and too large. In Figure 10 we compare the solutions at the final time: we can see that the NN based solution preserves the same symmetry properties of the exact solution, essentially non-oscillatory, and well captures the front of the wave. By analyzing the space-time plot of the viscosity introduced via the NN, as shown in Figure 11, we notice that the NN tends to add a larger amount of less localized viscosity when compared with EV for $k=1$. It is instead more similarly localized for $k=3,5$. More quantitative results about the error are shown in Table 2, we notice that the NN always outperforms the EV model in terms of $L^{1}$ error and overshoots/undershoots. It also obtains comparable results when considering mass variation (mv), an important metric to asses that the physics of the problem is respected. This comparison is limited by the choice of parameters $c_{K}$ and $c_{\max{}}$ we made and by the considered metrics. Indeed, while we manually tuned these parameters, there might exist a choice that leads to smaller error metrics for the EV model.
@@ 5.2.2 Test case 4: Burgers’ equation
In this test case, we keep the same parameters of the previous test case but we change the physical flux, namely, we employ a quadratic one, and the $\text{CFL}=0.15,0.4,0.4$. The Courant number is $C=0.3317,0.0924,0.0302$. These settings are more challenging due to the fact that as the solution evolves rarefaction fans and shock waves develop and collide. We compare the NN solution with a tuned EV model ($c_{K}=3.0,c_{\max{}}=1.0$) and two other not optimally tuned EV models. Figure 12 shows that one of the EV models ($c_{K}=1.0,c_{\max{}}=0.5$) is not able to add enough viscosity thus leading to large oscillation. The parameters used in this case are not far from the manually tuned: this showcases how time-consuming the parameter selection can be. In this case, the NN model is by far the best in avoiding overshoots/undershoots while correctly predicting the profile of the wave. The history of the viscosity in Figure 13 shows that as in the previous test cases, the NN tends to inject more viscosity than the EV model, in a more spread and less continuous way. For $k=5$ in particular, it is possible to see that the NN also tends to add dissipation in $x=0.2,t=0.35$ where the EV model did not. It remains to be investigated if this is a generalization error (since $k=5$ was not present in the training problems) or a feature of the discovered neural viscosity model. As shown in Table 3, the NN model outperforms all the EV models under all the considered metrics. This is an especially interesting result when considering that the reduction of the $L^{1}$ error for $k=3$ is of 25%, while also reducing overshoots/undershoots of almost 50%.
@@ 5.2.3 Test case 5: Sod shock tube
NN Ours EV NN [26] NN Ours EV NN [26] NN Ours EV
$\epsilon$ 5.7046e1 5.8845e1 5.9923e1 5.8787e1 7.2200e1 6.5057e1 4.3138e1 5.7406e1
$\nabla\epsilon$ 2.7103e1 2.7001e1 2.7639e1 1.3087e2 1.3356e2 1.3892e2 2.0463e2 2.1445e2
$[\![\epsilon]\!]$ 1.5025e0 1.8547e0 1.8261e0 1.0929e0 1.4114e0 2.3391e0 8.7019e-1 1.0149e0
o/u 3.0200e-1 7.5084e-1 6.3882e-1 2.0631e-1 7.5516e-1 1.7252e0 8.7122e-1 1.1302e0
mv 1.0951e-8 5.0988e-8 1.7928e-8 2.3772e-7 3.2462e-6 2.9315e-6 1.0951e-8 5.0988e-8
@@
It is endowed with constant Dirichlet boundary conditions (cf. [57]). For the discretization we select three cases with $k=1,3,5$, $h=\frac{1}{60},\frac{1}{30},\frac{1}{15}$, $\text{CFL}=0.27,0.61,0.88$, respectively. Even if this problem is present in the training dataset, the polynomial degree $k$ or CFL used to discretize it are different. The Courant number is $C=0.2970,0.1393,0.0785$. This is done to compare solutions with a similar number of degrees of freedom. The discontinuities in the initial conditions give rise to three characteristic waves, demanding robust numerical schemes to accurately compute the solution in the presence of shock waves and rarefactions. Namely, these are a right-moving contact wave and shock wave, and a left-moving rarefaction wave. Classical models aim to add a small amount of viscosity near the contact while constantly introducing dissipation in the region close to the shock. In Figure 14 we compare the NN solution at final time with the EV tune model with $c_{K}=1.0,c_{\max{}}=0.5$ and the NN method proposed in [16]. Our neural viscosity model provides a solution that has much smaller overshoots (see $x=0.2,k=1$ for instance) and that captures with more precision the wavefront (see $x=0.7,k=1$ for instance). In this case, it is particularly interesting to analyze the history of the viscosity shown in Figure 15. Namely, here we can notice a large discrepancy between what the NN and the EV model aim to do: the EV model stabilizes only the shock, meanwhile the NN identifies and tries to stabilize all three present waves. This is counterbalanced by injecting a smaller dissipation near the shock. For $k=5$, our model injects a large amount of viscosity near the contact discontinuity, which does not make physical sense. However, it is hard to identify the precise reasons why the model introduces such a large amount of artificial viscosity. Indeed, one of the limitations of our hybrid approach is that the model learns the viscosity from the data without explicit guidance on what constitutes correct or incorrect behavior in specific contexts. Despite being more intrusive, our NN outperforms the EV model and the NN of [16] under all the considered metrics, as shown in Table 4, where we report the error metrics in time and their cumulative value.
@@ 5.2.4 Test case 6: Shu-Osher problem
NN Ours EV NN [26] NN Ours EV NN [26] NN Ours EV
$\epsilon$ 2.6293e3 3.0574e3 3.6781e3 2.2953e3 4.0510e3 3.1186e3 3.7385e3 5.5862e3
$\nabla\epsilon$ 1.7160e3 1.8731e3 2.0671e3 5.6320e3 7.8568e3 6.1255e3 1.3576e4 1.6539e4
$[\![\epsilon]\!]$ 1.2497e2 1.0530e2 8.9058e1 9.8602e1 7.1604e1 7.2939e1 5.2003e1 5.4259e1
o/u 3.1512e0 3.9106e0 1.0716e0 2.8027e0 5.7163e0 2.5512e0 3.8023e0 4.6424e0
mv 4.2955e-7 2.2112e-6 1.0787e-6 8.0499e-5 2.2317e-4 1.1017e-4 7.3125e-7 4.5271e-6
@@
The problem is completed with Dirichlet boundary conditions on the left and Neumann on the right part of the boundary (cfr. [56]). Since the solution exhibits a high frequency wave we employ a finer discretization, namely $h=\frac{1}{150},\frac{1}{75},\frac{1}{50}$ with $k=1,3,5$, and $\text{CFL}=0.12,0.3,0.4$, respectively. The Courant number is $C=0.2707,0.1337,0.0762$. From Figure 16, we can see that for the case $k=1$ the NN and the EV model and the NN of [16] produce a rather similar solution. However, the quantitative analysis reported in Table 5 shows that the neural viscosity model is slightly more accurate in the $L^{1}$ norm, and when considering overshoots and undershoots. Instead, the cases $k=3,5$ show that the EV model and the NN of [16] smooths the solution too much, and thus it is not able to capture the high frequencies. On the other hand, our NN shows much better results, namely the solution correctly represents the high-frequency wave pack present between $x=0.5$ and $x=2.5$. This also proves that the neural viscosity model is able to exploit the high-order nature of the method, indeed the three considered cases have the same number of degrees of freedom. In particular, for $k=3$ we have a reduction of the $L^{1}$ error of 43% and of 51% of overshoot/undershoot. The dissipation injected by the NN, as shown in Figure 17, is concentrated on the shock and, differently from the EV model, adds also other small patches of viscosity at the boundary of the sinusoidal wave pack.
@@ 5.3.1 Test case 7: linear advection
Moreover, it gives us the possibility to check the mass conservation properties of the neural viscosity model. We solve the problem in $\Omega=(0,2)^{2}$ with a final time $T=0.25$ on a structured grid with $7200,1800,450$ triangular elements and $k=1,3,5$, $\text{CFL}=0.075,0.3,0.32$, respectively. The Courant number is $C=0.1061,0.0471,0.0181$. We focus on $k=1$ since the other two cases are analogous. As done in the previous test cases, we compare the neural viscosity model with a tuned EV model ($c_{K}=1.0,c_{\max{}}=0.5$) and two other EV models that inject a sub-optimal amount of dissipation (too large $c_{K}=2.0,c_{\max{}}=1.0$ and too small $c_{K}=0.5,c_{\max{}}=0.25$). From Figure 18 we clearly see that both the NN and the EV model smooth the discontinuous profile of the initial condition in order to avoid oscillations. The NN solution is symmetric and visually exhibits a sharper wavefront. This is confirmed by the quantitative analysis of the error reported in Table 6. In particular, it is interesting to observe that the NN model injects more viscosity than a classical EV model but has a smaller $L^{1}$ error. This is not an intuitive behavior when comes to the EV model since from both Figure 19 and Table 6 it is clearly observable that EV models that inject larger dissipation have smaller overshoots/undershoots but larger $L^{1}$ error. Moreover, the solution obtained with the NN conserves the mass better than the one obtained with the EV model, showing that also physics is respected. Therefore, in this case, the neural viscosity model is effectively learning a new model that has better properties than the EV one.
@@ 5.3.2 Test case 8: KPP rotating wave
NN Ours EV NN [26] NN EV NN [26] NN EV
$\epsilon$ 1.6196e5 3.9369e5 6.3515e5 1.6956e5 3.8346e5 5.6881e5 2.5583e5 2.8887e5
$\nabla\epsilon$ 2.4551e5 4.3755e5 5.4866e5 1.2625e6 1.7127e6 1.8699e6 2.6152e6 2.8368e6
$[\![\epsilon]\!]$ 1.0973e5 1.4006e5 9.5575e4 8.2736e4 9.7502e4 6.4618e4 4.4259e4 6.2624e4
o/u 8.0091e3 1.9289e4 4.8742e3 6.4779e3 6.8667e3 2.5480e3 4.6155e3 5.0963e3
@@
The problem is endowed with periodic boundary conditions. For discretization, we employ a structured triangular mesh with $7200,1800,450$ elements and $k=1,3,5$, $\text{CFL}=0.11,0.4,0.48$, respectively. The Courant number is $C=0.11,0.0444,0.0192$. We compare the neural viscosity model with EV models tuned like in the previous test case and the NN of [16]. The results reported in Figure 20 and show that the neural viscosity model produces more accurate solutions with respect to the EV one. In particular, it presents less wriggles and a sharper wave front. To understand the behavior of the model we report also in Figure 22 the maximum, mean, and median of the injected viscosity. As was the case in 1D, the NN model injects more viscosity than the tuned EV model, but retains a lower $L^{1}$ error, as shown in Table 7. On the other hand, the method proposed in [16], is the one that produces the smoothest solution, featuring far less oscillations than the other two methods. However, it is also the less accurate solution from the point of view of the $L^{1}$ norm.